Database last updated on Fri, 30 Jan 2026 13:41:07 +0000 by localhost

Critical warnings from b-files' comments for days between 2024270 and 2024270 (0 sec, 5 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
rl:2024 270 19 06 42 Motor #3: appears stuck with a reference sensor activated.31B27024.224TORON#22405:13:561-1
rl:2024 270 19 29 19 New value set for watermark SUPPLY.VOLTAGE.HIGH303B27024.224TORON#22405:14:071-1
hg:Lamp not on... hg test terminated F1= 0 F4= 11861B27024.224TORON#22407:04:151-1
hg:Lamp not detected f1= 0 f4= 11861B27024.224TORON#22407:04:161-1
pd:Command file pd not found --- continuing11B27024.224TORON#22418:00:291-1

Instrument constant changes for the week before 2024270 (1 sec, 0 rows)

Hg with step corrections larger than 10 steps or multiple corrections larger than 2.5 steps (0 sec, 1 rows)

SerialYYYYJJJHG largemax correction
22420242705118.9

HP with step corrections larger than 20 steps (0 sec, 1 rows)

SerialYYYYJJJHP largemax correction
22420242701280.0

Non-critical warnings from b-files' comments for days between 2024270 and 2024270 (0 sec, 11 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
rl:2024 270 19 06 41 Reset No -> Tepid from 2024 270 18 53 47 until 2085 234 12 38 20 . Counts: 1,013B27024.224TORON#22405:13:521-1
rl:2024 270 19 06 42 Checksum bad in low level command.13B27024.224TORON#22405:13:541-1
rl:2024 270 19 06 42 Motor #3: appears stuck with a reference sensor activated.13B27024.224TORON#22405:13:561-1
rl:2024 270 19 06 42 Motor #11: appears stuck with a reference sensor activated.13B27024.224TORON#22405:13:591-1
rl:2024 270 19 07 40 Reset Warm -> Warm from 2024 270 19 07 37 until 2024 270 19 07 38 . Counts: 1,113B27024.224TORON#22405:14:021-1
dp:Doing dispersion with internal HG.199B27024.224TORON#22410:19:351-1
pd:Command file pd not found --- continuing199B27024.224TORON#22418:00:29continuing1-1
uv:UV extended scan aborted23B27024.224TORON#22418:52:231-1
re:Azimuth discrepancy = 7948199B27024.224TORON#22419:09:037948-3030
ds:DS intensity too high ( 325765/cy) quitting.199B27024.224TORON#22420:23:561-1
ds:DS intensity too low ( 45/cy), skipping.399B27024.224TORON#22422:27:341-1

Measurements and tests summary.


_jday_loc_serial_skcdsdtedhghprssluvzs
270TORONTO#2240 0 2967.28 0.0 1867 29,inside,menu,Must do HG now,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.46499),Saving backup OP,Starting schedule ecsc,Waiting until ZA = -100 now -101.1,Waiting until ZA = -80 now -80.9,Waiting until ZA = -85 now -85.6,Waiting until ZA = -85 now -87.3,Waiting until ZA = -85 now -89,Waiting until ZA = -85 now -91.2,Waiting until ZA = -85 now -9319725275610117
(1 sec)

Number of days since last dome/window cleaning (0 sec, 2 rows)

serialdays
224more than a year

Missing files (0 sec, 1 rows)

Seriallast datadays ago
22420242710-1

Run/Stop results for last week. (0 sec, 1 rows)

SerialTotalFailed 0Failed darkFailed 1Failed 2Failed 3Failed 4Failed 5Failed double
2241701000000

AP results for SB Brewers. (0 sec, 1 rows)

SerialJdayt1 offt1 ont2 offt2 ont3 offt3 onHV offHV on+12V off+12V on+5V off+5V on-12V off-12V on+24V off+24V on+5V sps off+5V sps onSL A offSL A onSL V offSL V onWatts
224202427027.427.428.628.627.527.51275.11273.912.012.05.25.1-12.1-12.124.124.15.15.10.01.50.010.4615.8

Standard lamp intensity analysis (0 sec, 1 rows)

Serialminmaxfilterfixable30-day history
#224267222026853800yes